# LLM Trading: Analysis of LLM Agent Behavior in Experimental Asset Markets

### v1

@@ 5.2 Market Analysis
MSE (Fundamental): 5.8583  PCC (Human Avg): 0.2673
Model-Level Observations
As shown in Table 2, the final portfolio values across agent types are relatively close together, with no single participant dominating the market. In each Battle Royale market, the spread between the highest and lowest portfolios is fairly moderate—indicating that while some agents capitalize on slightly better trading opportunities, the market does not produce extreme winners or losers. This relatively tight distribution suggests that, in the mixed LLM-human environment, most participants converge on trading strategies that yield comparable overall returns. Consequently, although small differences do emerge, the competitive pressure in these mixed markets appears to limit the degree of inequality in final outcomes.
@@ 6 Discussion and Conclusions
In this study, we investigated how Large Language Models (LLMs) behave in an experimental finance paradigm that has historically elicited asset bubbles and crashes when conducted with human participants. Through both single-model markets and mixed-model “battle royale” competitions, we found that LLM-based agents typically converge on prices near the fundamental value, thereby exhibiting more “textbook-rational” behavior than human subjects. Although some single-model markets did display mild bubble-like phenomena—particularly GPT-3.5, which inflated prices without a clear crash, and Mistral and Grok 2, which produced modest run-ups, —these tendencies were weaker than what is commonly observed in human-only markets. Moreover, individual outcomes in LLM-driven markets tended to cluster more tightly around similar final portfolio values, indicating less strategic heterogeneity than seen among humans. These findings collectively suggest that although LLMs may serve as rational traders, they seldom replicate the pronounced boom-and-bust patterns that emerge from human behavioral biases. Therefore, our results caution against using LLMs as replacements or for piloting for human participants in experimental finance studies. Future work may explore whether new prompting strategies, incentives can induce more human-like trading patterns from LLM-based agents, but for now it remains clear that these models trade with a consistency and rationality that diverge significantly from the behaviors displayed by human counterparts.
Acknowledgments
The authors extend their sincere gratitude to Geeling Chau, Pantelis Vafidis, and Dr. Andrew Sinclair for their invaluable advice and guidance throughout this project. Their insightful feedback and generous support were instrumental in shaping the direction and scope of this work.
@@ Appendix A Additional Experiment Details
Below, we describe in greater detail the sequence of screens shown in the human-version of the experiment. Note that (4) is omitted in our adaptation of the experiment for LLM-based agents and screens (1) - (3) are condensed into a single ”screen”, or more accurately call to the model. Please refer to Appendix B for further discussion of the adaptation of the experiment.
@@
    Order Submission Screen (20 seconds): Participants submit orders to buy or sell the asset. To simplify the interface, all orders are entered via a grid with quantity along the horizontal axis and price along the vertical axis. Prices are restricted to be within ±$\pm 20\%$ of the current market price.
@@
    Forecast Screen (30 seconds): Participants submit their expectations about current and future market prices. Price forecasts are made by positioning a slider. All sliders had a lower limit of zero, while the upper limits varied depending on how far out the forecast was made. For the current round and two rounds hence, the upper limit was twice the current market price. The five-round forecast was limited to 2.5 times the market price, and for the ten-round forecast, the limit is set to triple the market price. To aid in the aesthetics of the sliders, the upper limit was rounded up to the nearest tick mark. For example, if the current market price is 36, the upper limit for the current round forecast is 72. If the tick marks on the slider are determined to be spaced at 5, the upper limit is rounded up to 75.
@@
    Round Results Screen (10 seconds): Participants see the actual realized price for that period, how many shares were traded in total, and how many shares they traded.
  4. 4.
    Risk-Elicitation Screen (20 seconds): Participants see 4 pairs of lotteries that change as they select riskier or less risky lotteries. This task applies the Dynamically Optimized Sequential Experiment (DOSE) to the Holt & Laury (2002) 10-question set, as outlined by Chapman et al. (2024)
Figure 3: Example task screens shown to human subjects
At the conclusion of the experiment, the risky asset is redeemed for a fixed value of 14 units, representing the fundamental price, calculated using the formula presented in Equation 1. Note our experiment paradigm was designed such that this fundamental value remains constant through the entirety of experiment. This enables us to objectively classify when the market is in a ”bubble” state, given we always have a constant fundamental value to compare market price to.
@@ A.1 Price Formation
Cumulative sell quantity is likewise defined as:

### v3

@@ 5.1 Experimental Setup
MSE (Fundamental): 5.86  PCC (Human Avg): 0.27
5.2 Battle Royale Market Analysis
Across three heterogeneous markets, two converged on prices near fundamentals, while one produced a modest bubble resembling human markets but with a lower peak. Trading volume spiked at the end of this bubble run, echoing the late-round rush seen in human experiments (Figure 5.1). Overall, these dynamics align most closely with the Rational (R) hypothesis, although they demonstrate that competitive interaction can occasionally generate bubble-like patterns.
At the agent level (Table 2), portfolio values cluster tightly, with no model consistently dominating. The spread between best and worst traders remains small, suggesting convergence on broadly similar strategies and limited inequality of outcomes in mixed markets. This shows that none of these models are able to consistently best the others in competitive trading.
6 Trading Strategy Analysis
After each round, LLM agents were prompted to provide details of the strategy they used to drive their trading behavior. This was done both to allow us to analyze their behavior more deeply and to enable the LLMs to develop long-running strategies, utilizing Chain-of-Thought-like reasoning. In contrast, human subjects were asked what approach they employed only at the end of the experiment. More details of our data cleaning techniques and additional analysis can be found in Appendix F.
We first examined the word frequency of both humans and LLMs in the mono-agent runs. Of the top six words for each group, five are the same, though in a different order (”sell”, ”buy”, ”stock”, ”price”, and ”market”; see Table 6 in Appendix for details). The top ten most popular words for the human subjects include ”low” and ”high”, which do not appear among the top ten words used by the LLMs. Conversely, the words ”continue” and ”buyback” are among the most common for the LLMs (see also word clouds in Figure 22 in the Appendix). This pattern suggests that human traders follow a ”buy low, sell high” strategy in an attempt to outperform the market. In contrast, LLM agents adopt a longer-term approach that emphasizes the asset’s intrinsic value.
6.1 Text Mining and Strategy Comparison to Humans
To assess whether LLMs engaged in coordination or converged on shared narratives, we ran a text–mining analysis of the “insight” and “plan” files. We found no explicit coordination language (0 hits on terms such as ”coordinate” or ”work together”), and only a handful of generic references to ”anticipating other agents.” By contrast, we observed several instances of near–orthogonal language between agent pairs (cosine similarity $\leq 0.1$; e.g., “buy aggressively” vs. “sell off”), suggesting divergence rather than cooperation. Overall, mixed–model markets appear to consist of independent traders reacting to shared price signals rather than colluding.
We then directly compared LLM and human strategies. A classifier trained on the insights/plans text achieved >>99% accuracy in identifying which model produced which text. Extending this to a seven–way classifier including humans (using their end of experiment reflections), we found that $\sim$33% of Gemini-1,5 Pro and GPT-4 reflections were misclassified as human, while all other models were below 10%. Binary classifiers confirmed this pattern: GPT-4 text was misclassified as human nearly 50% of the time (Mistral-Large 36%; GPT-3.5 28%), while other models were below 3%. Despite superficial stylistic overlap, cluster analysis of strategy words showed almost no clusters containing both humans and LLMs (in one run, only 8/497 humans clustered with any model), suggesting that the underlying reasoning remains distinct.
To probe this distinction further, we ran a Latent Dirichlet Allocation (LDA) Blei et al. (2003) on all strategies combined. Using the Python gensim library Řehůřek & Sojka (2010), we assigned each strategy to one of two topics. The first topic, identified by the words orders, sell, buy, shares, stock, denoted strategies aligned with a “buy low; sell high” practice. The second topic, identified by the words stock, adjust, buyback, closely, current, reflected a more fundamental value focused approach. The LDA assigned 89.3% of human strategies to the first topic but only 36.6% of the LLM strategies; conversely, 63.4% of LLM strategies (10.7% of humans) fell into the second category. A t-test confirmed this difference is highly significant ($p<.001$). Figure 23 shows this distribution.
6.2 Linguistic Audit of Model Strategies
We also wanted to identify why specific models generate larger, more human-like bubbles, while others adhere relatively closely to fundamentals. To do so, we ran a finer-grained linguistic audit of the ”insight” and ”plan” outputs. Each statement was tagged as speculative, fundamental, or generic based on keyword dictionaries.
Table 3: Distribution (%) of speculative, fundamental, and generic text in agent strategy statements. The behavior column reflects observed market outcomes.
Model Speculative Fundamental Generic Behavior
Grok-2 68.4 16.0 15.6 Bubble-prone
GPT-3.5 28.4 18.0 53.6 Bubble-prone (No Crash)
GPT-4o 17.9 59.6 22.5 Bubble-dampening
Claude-3.5 Sonnet 14.5 82.8 2.7 Bubble-dampening
Gemini-1.5 Pro 20.5 58.7 20.8 Bubble-dampening
Mistral-Large 35.8 52.0 12.2 Mixed
Speculative keywords: ”trend”, ”momentum”, ”price rising”, ”price falling”, ”buy more”, ”bubble”, ”sell off”, ”rally”.
Fundamental keywords: ”undervalued”, ”intrinsic”, ”dividend”, ”fair value”, ”discount”, ”fundamental”, ”14”, ”buy back”.
Grok-2 and GPT-3.5 seem to create bubbles because speculative narratives dominate their reasoning, whereas GPT-4o, Claude-3.5, and Gemini-1.5 Pro seem to dampen exuberance by emphasizing intrinsic value anchors. Mistral-Large oscillates between the two, producing smaller, transient bubbles with a trajectory of value $\rightarrow$ momentum $\rightarrow$ uncertainty. These linguistic signatures provide a possible mechanistic explanation for the divergent market outcomes observed across models.
7 Market Treatments
Beyond the baseline constant–fundamental design, we tested whether LLM’s rationality profiles held under more complex and dynamic market conditions, such as when payoffs and fundamental values shift over time.
To test whether our findings extend beyond constant fundamental value, we conducted two separate mono-agent market experiments (each with three sessions) featuring a mid-experiment dividend shock. Beginning in Round 15, we either doubled (7.1) or halved (7.2) the dividend and redemption value, which doubles/halves the fundamental value of the asset. The agents were not informed of this change prior to the experiment and were told at Round 15 that the dividend and redemption value had changed (either doubled or halved) and would remain so for the rest of the experiment. This mirrors classic shock tests in experimental finance, and simulates how unexpected ”news” can impact the value of an asset, allowing us to assess whether LLMs adjust appropriately when the fundamental value changes. Note we were unable to run Grok 2 as the public endpoint has been deprecated.
7.1 Dividend (”News”) Shock Doubling
Results show that all models except GPT-3.5 quickly converged toward the new fundamental value. Claude-3.5 Sonnet, GPT-4o, Gemini-1.5 Pro, and Mistral-Large all incorporated the new information, though the process was not instantaneous, stabilizing prices near \$28. GPT-3.5, however, lagged behind: its forecasts and orders remained anchored near the pre-shock equilibrium. Notably, Gemini-1.5 Pro occasionally overshoots the FV, creating minor bubbles and crashes.
Figure 3: Mono-Agent Market Price Charts Dividend Shock Doubling
Models slowly adjust to the new FV price, with the exception of GPT-3.5. We observe that Gemini-1.5 Pro occasionally creates a modest bubble.
7.2 Dividend (”News”) Shock Halving
As shown in Figure 4, all models except GPT-3.5 converged toward the new fundamental value. Claude-3.5 Sonnet, Gemini-1.5 Pro, and Mistral-Large all incorporated the new information within 2–3 rounds, stabilizing prices near \$7. GPT-3.5 and GPT-4o, however, lagged behind, with GPT-3.5’s prices remaining anchored near the pre-shock equilibrium, and GPT-4o incorporating the information slowly.
Figure 4: Mono-Agent Market Price Charts Dividend Shock Halving
All of the markets, with the exception of GPT-3.5, quickly adjust to the new fundamental value after the dividend shock. We observe that GPT4 operates within a modest bubble before converging to its redemption value towards the end of the market session.
These treatments reinforce our main result: most LLMs exhibit a fundamental value oriented adjustment, while GPT-3.5 is uniquely prone to inertia and bubble-like dynamics. We notice a slight increase in the propensity for the models to create bubbles, but they still occur infrequently. It may appear that the models converge more quickly to downward shocks than upward shocks. Please keep in mind that upward shocks are larger in absolute magnitude, and our experiment design limits orders to $\pm$ \$3 of the current price. This was done to make the experiment accessible to humans and carried over to our LLM experiments for parity.
We also conduct multiple robustness checks in the Appendix, ensuring that slight experimental differences (such as removing the risk-elicitation task for LLMs D.2) or LLM trading experience (D.1) do not materially alter our conclusion. While these treatments represent only a small subset of the numerous possible experimental modifications, they demonstrate that our conclusion regarding LLM rationality remains robust to various changes in market conditions.
8 Rationality Properties of Individual LLM Forecasts
In addition to analyzing the pricing behavior of LLMs, we were also interested in how these models generate forecasts for future price periods. We captured each model’s forecasts, which allowed us to assess not only the rationality of their forecasts, whether they align with fundamental valuation principles, but also whether the LLMs themselves behave rationally in response to their expectations. We compare this analysis to that of our human traders to understand whether LLMs, in the mono-agent market setting, make similar or different forecasting mistakes compared to their human counterparts. The mathematical formulation for calculating the mean and median forecasts for any given session is provided in the Appendix (Section G.3).
Forecast errors are defined as $E_{i,t}^{h}=P_{t+h}-f_{i,t}^{\,t+h}$, where $P_{t+h}$ denotes the actual price at round $t+h$ and $f_{i,t}^{\,t+h}$ is agent $i$’s forecast formed at round $t$ for horizon $h$; positive values indicate underprediction. Full distributions and horizon-specific test results are reported in Appendix G.3. Humans exhibit substantial systematic underprediction: the mean one-step error ($h=1$) is $1.67$. By contrast, all LLMs’ mean errors are close to zero. Consistent with this, Appendix Figure 26(b) shows that LLM forecast-error distributions are more tightly centered around zero than those from human markets, indicating higher forecast precision.
Figure 5: Mean Forecast Error by Round
Humans exhibit substantial systematic underprediction, with a mean one–step forecast error of $1.67$. By contrast, all LLMs are much closer to zero.
We also applied three additional diagnostics to compare the forecast-rationality profiles of LLM agents and human subjects: (i) unbiasedness—whether forecast errors are centered on zero (no systematic bias); (ii) zero autocorrelation—whether errors in one round fail to predict errors in later rounds; and (iii) errors uncorrelated with forecasts—whether the forecast itself fails to predict the size of the error (no optimism/pessimism bias).
Results are summarized in Table 4 and expanded upon in Section G.3 of the Appendix. Results suggest that LLMs generally forecast more accurately than humans in the short run, but their rationality profiles differ between models. GPT-3.5 is highly biased and correlated; Claude-3.5 and GPT-4o are the most consistently rational; Gemini-1.5 Pro, Grok-2, and Mistral-Large fall in between. Importantly, no model fully reproduces the human error pattern.
Table 4: Forecast Accuracy and Rationality Properties of Humans and LLMs.
Model Forecast Error Unbiasedness Zero Autocorr. Errors $\perp$ Forecasts
Humans 1.67 0.56 0.17 0.31
Claude-3.5 Sonnet 0.09 0.46 0.08 0.16
Gemini-1.5 Pro 0.14 0.41 0.18 0.41
GPT-3.5 –1.54 0.00 0.60 0.00
GPT-4o 0.04 0.40 0.60 0.03
Grok-2 –0.31 0.56 0.18 0.27
Mistral-Large 0.00 0.44 0.14 0.36
Notes: ”Forecast Error” is the mean one–step error ($E_{i,t}^{0}$; positive = underprediction). Other columns report the proportion of agents passing each rationality test, averaged across horizons $h\in\{0,2,5,10\}$. Full horizon-specific results are reported in the Appendix.
9 Potential Limitations and Broader Impacts
Our results should be interpreted in light of several caveats. First, LLM agents do not experience incentives, emotions, or online learning in the human sense; their “strategies” emerge from the prompt structure and frozen weights, rather than endogenous reasoning or reinforcement. Second, we study stylized markets with a single asset and a short horizon. How LLM traders respond to longer horizons, short selling, or richer strategic interactions remains an open question. Finally, we evaluated only a handful of foundation models without fine-tuning. Model choice, prompting tactics, or retrieval-augmented memory may materially alter behavior.
Despite these limitations, our work has three areas of immediate impact. First, in live markets, several hedge funds have begun deploying LLM-driven order flow. As machine agents come to represent a measurable share of trading volume, both regulators and practitioners will need empirical baselines to understand the dynamics they generate. Second, almost all significant findings in behavioral finance (loss aversion, herding, mental accounting) were initially discovered through lab experiments (Kahneman & Tversky (2013), Thaler (1985)). As we attempt to understand how LLMs trade and the biases they exhibit, we thought it best to adopt the same approach. Third, for experimental design, recent papers have explored the idea of substituting LLMs for human subjects in laboratory finance experiments (Cui et al. ); our evidence of systematic departures from human behavior underscores the risk of doing so without rigorous validation. Therefore, we advise caution in replacing human subjects with ”off-the-shelf” LLM agents. We acknowledge these limitations and thus view this study as a starting point, rather than an endpoint, for incorporating LLM agents into financial and behavioral research.
@@ 10 Discussion and Conclusions
In this study, we compare the behavior of Human and Large Language Models (LLMs) in an experimental finance paradigm that has historically elicited asset bubbles and crashes when conducted with human participants. We found that LLM-agents typically converge on prices near the fundamental value, thereby exhibiting more “textbook-rational” behavior than human subjects. Although some single-model markets did display mild bubble-like phenomena, particularly GPT-3.5, which inflated prices without a clear crash, and Mistral-Large and Grok 2, which produced modest run-ups, these tendencies were weaker than what is commonly observed in human-only markets. Moreover, individual outcomes in LLM-driven markets tended to cluster more tightly around similar final portfolio values, indicating less strategic heterogeneity than seen among humans. We found that these findings persisted even under changing market conditions, news/dividend shocks. These findings collectively suggest that although LLMs may serve as rational traders, they seldom replicate the pronounced boom-and-bust patterns that emerge from human behavioral biases. We also analyze the forecasting behavior of these LLM agents and find that the majority of the LLMs generate more accurate forecasts of future price behavior than human traders; however, we observe some heterogeneity among LLM agent models in their ability to forecast rationally. Across the board, the LLM agents performed differently from humans on most tasks; therefore, our results caution against using LLMs as replacements or for piloting for human participants in experimental finance studies. Future work may explore whether new prompting strategies or incentives can induce more human-like trading patterns from LLM-agents. However, for now, it remains clear that these models trade with a consistency and rationality that diverge significantly from the behaviors displayed by human counterparts.
11 Ethics Statement
We do not believe there are any ethical concerns with this work, all human data was collected under review of the IRB at [University]. The experiments with large language models involved only synthetic agents and did not expose participants or sensitive populations to potential harm. All analyses follow standard practices in experimental economics and machine learning. We believe this work adheres fully to the ICLR Code of Ethics and raises no issues regarding privacy, fairness, security, or potential misuse.
12 Reproducibility Statement
All code will be released alongside the camera-ready submission. Complete experimental details, including instructions provided to agents and human participants, as well as the full set of prompts, are documented in Appendix A and Appendix I. These materials, together with the statistical procedures and robustness checks reported in Section 4 and Appendices E–G, ensure that all results can be readily reproduced. Our goal is not only to enable replication of the experiments presented here, but also to provide a foundation that future work can extend and build upon.
13 LLM Disclosure
Large Language Models (LLMs) were used in a limited capacity to assist with LaTeX formatting and debugging. They were not used for research ideation, experimental design, data analysis, or substantive writing. The authors take full responsibility for the content of this paper.
@@ Appendix A Additional Experiment Details
Our experimental design involves an asset market where subjects trade a risky asset in an open-call market for 30 periods (Smith et al., 2014; Holt et al., 2017). Participants are given 4 shares of stock and 140 units of cash. Participants have two options: either invest cash in the risky asset, which pays out dividends of either 0.4 or 1 with equal probability, or hold cash, which accrues interest at a rate of 5% each trading period. The dividend payment is drawn independently for each round.
The experiment consists of a sequence of four screens in this order, as depicted in the Figures below:
@@
    Order Submission Screen (20 seconds)(Figure 6): Participants submit orders to buy or sell the asset
    Figure 6: Order Submission Page
@@
    Forecast Screen (30 seconds)(Figure 7): Participants submit their expectations regarding current and future market prices.
    Figure 7: Forecast Page
    Participants are asked to submit their estimation for the current market period and periods 2, 5, and 10 rounds ahead.
@@
    Round Results Screen (10 seconds)(Figure 8): Participants see the actual realized price for that period, how many shares were traded in total, and how many shares they traded.
    Figure 8: Round Results Page
After the experiment, the risky asset is redeemed for a fixed value of 14 units, representing the fundamental price, calculated using the formula presented in Equation 7. Note that our experiment paradigm was designed such that this fundamental value remains constant throughout the entirety of the experiment. This enables us to objectively classify when the market is in a “bubble” state, given we always have a constant fundamental value to compare the market price to.
@@ A.1 Price Formation
The cumulative sell quantity is likewise defined as: